OpenStudy (anonymous):

need help with this trash.. (lol) A neighborhood has provided trash cans to residents to dispose of their waste. All trash cans have a diameter of 0.5 meters and a height of 1.5 meters. What is the minimum number of trash cans required to dispose of 10 cubic meters of waste? Answer 34 102 33 101

OpenStudy (oaktree):

Volume of one can: pir^2h = pi x (1/16) x 3/2 = 3pi/32 = 0.2945 That goes into 10 cubic meters how many times? 10 / (3pi/32) = (10 x 32) / 3pi = 33.9 = 34 cans

OpenStudy (anonymous):

ah ok, ty I get it. Idk why i wasn't getting 0.2945 b4..